OpenStudy (anonymous):

what is sin squareroot7/4?

OpenStudy (anonymous):

Calculator can do that

OpenStudy (anonymous):

i need to find cos

OpenStudy (jhannybean):

\[\large \sin(\sqrt{\frac{7}{4}}) = \sin (\frac{\sqrt{7}}{2})\] one way you could find cos is by squaring sin, and using the identity sin^2 (x) + cos^2 (x) = 1.

OpenStudy (jhannybean):

Err...i mean you can take the derivative of sin(sqrt (7)/2)

OpenStudy (anonymous):

i can't figure out the answer

OpenStudy (nirmalnema):

sin theta = p/h cos theta=b/h =\[\sqrt{45}/\sqrt{7}\]